Jane can complete a sales route by herself in 45 minutes. Working with an assistant she can complete a sales route in 40 minutes. How long would it take her assistant to do it alone?
Answer by jorel1380(3719)

Let n be the amount of time it takes for the assistant working alone. Then:
1/45+1/n=1/40
8n+360=9n
n=360
It would take the assistant 360 minutes, or 6 hours, to do the route alone. ☺☺☺☺